OpenStudy (anonymous):

can someone help check my algebra 2 work plz??

OpenStudy (anonymous):

the question is You must: Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships WARNING! No identities used in the lesson may be submitted. Create your own. See what happens when different binomials or trinomials are combined. Below is a list of some sample factors you may use to help develop your own identity. (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d)

OpenStudy (anonymous):

my work is (x-y)(x+y)=x^2-y^2 (x-y)(x+y)=x^2-xy+xy_y^2 then you simplify it to x^2-y^2 which proves (x-y)(x+y)=x^2-y^2

OpenStudy (anonymous):

is that right please help

OpenStudy (anonymous):

Yes. That's right. I would start from line two and then your last line.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

(x-y)(x+y) = x^2 -xy +xy -y^2 = x^2 - y^2, thus (x-y)(x+y) = x^2 - y^2

OpenStudy (anonymous):

oh ok thx so how do i Demonstrate that my polynomial identity works on numerical relationships

OpenStudy (anonymous):

Pick some numbers, i'm guessing? Say, let x = 2 and y = -1, show identity holds.

OpenStudy (anonymous):

so just plug in random numbers??

OpenStudy (anonymous):

That's what I would do.

OpenStudy (anonymous):

ok thank you so muchhhhhhhhhh

OpenStudy (anonymous):

so i would plug in x=2 y=-1 and get

OpenStudy (anonymous):

im plugging it into both sides right??